# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v2

@@ 9.3 On Research Question 3
Recent research trends in the prediction domain focus on enhancing the robustness and generalization of prediction models. As discussed in Section 5, a significant research gap exists in quantifying and mitigating uncertainty within prediction pipelines. One approach to mitigate aleatoric uncertainty is through collaborative perception, presented in Section 8, which compensates for individual perception system limitations through collaborative holistic view of the environment. Planning-conditioned trajectory prediction methods, presented in Section 6, represent a promising direction by incorporating motion planning principles into prediction frameworks. By modeling planned trajectories and potential destinations, planning-conditioned approaches provide interpretable and constraint-compliant predictions, particularly in interactive and complex scenarios.
@@ 9.4 Future Directions
    As noted in subsection 2.2, in the detect-track-predict paradigm, the input to the prediction pipeline consists of tracklets that inherently carry detection and tracking errors. However, most existing methods assume error-free past trajectories during inference, which is an unrealistic assumption. A key research gap lies in evaluating the robustness of prediction models under perception noise, and in designing tolerance mechanisms or training strategies that allow models trained on clean data to generalize to noisy or incomplete inputs.
  4. 4.
@@
    Handling out-of-distribution scenarios remains an active research area. One promising direction involves a dual-process strategy: using predefined, optimized algorithms in familiar scenarios while leveraging the generalized reasoning capabilities of Vision-Language Models (VLMs) in unfamiliar or uncertain situations. Another approach focuses on simulation-to-reality (Sim2Real) adaptation. Sim2Real techniques involve generating synthetic scenes that mimic real-world conditions, enabling models to learn in controllable environments where varying levels of diversity can be introduced before deployment. Additionally, continual learning during trial deployments or pre-deployment phases can help tailor models to local conditions. A small dataset specific to the target city may be collected to fine-tune the model while preserving previously learned knowledge. Although some works, such as [654, 655, 656], explore this direction, further research is needed to fully realize its potential in autonomous systems.
  6. 6.
@@
    Autonomous vehicles must predict trajectories for diverse agent categories, including vehicles of various sizes (e.g., trucks, vans, sports cars), pedestrians, bicycles, and motorcycles. Each agent type exhibits distinct kinematic properties, typical speeds, and acceleration patterns, which significantly influence their motion behavior. Vehicle trajectories are constrained by factors such as steering limits and acceleration bounds, whereas pedestrian motion is more unconstrained and inherently stochastic. A key challenge is developing a unified prediction framework capable of accommodating these heterogeneous characteristics. One approach involves using a single model trained on large, diverse datasets, with agent category labels and relevant priors provided as inputs, or following a hierarchical design where the model first classifies the agent type and then delegates prediction to category-specific modules. This setup requires carefully curated training data to ensure balanced learning across categories. Alternatively, using category-specific predictors can improve specialization but increases inference time and system complexity.
  7. 7.
@@
    Trajectory prediction models find integration in both fully autonomous vehicles and modern advanced driver-assistance systems (ADAS), contributing significantly to road safety. In full-stack autonomy, prediction modules inform planning about potential collisions. Automotive companies like Waymo and Cruise integrate multi-agent forecasting in their planning pipelines [^9] https://waymo.com/research/motionlm/. In ADAS [^10] https://www.mobileye.com/blog/ prediction models help infer lane-change intent and pedestrian movement, enabling early warnings that give human drivers more time to maneuver and avoid collisions. Another application involves integrating prediction with emergency braking systems, as seen in Mercedes-Benz vehicles, which use predictive cues to autonomously stop when drivers respond too late [^11] https://group.mercedes-benz.com/innovation/product-innovation/technology/pedestrian-emergency-braking-system.html. Future research should continue on integrating prediction models in traffic management systems to enable proactive safety interventions.
@@ 10 Conclusion
Beyond the taxonomy and evaluation, this study highlights the pressing challenges and evolving trends in the field. Sections on robustness and generalizability emphasize the need for uncertainty quantification and mitigation, while collaborative perception is discussed as a promising solution for addressing occlusions, sensor noise, and restricted sensing capabilities. The review also explores planning-conditioned trajectory prediction, which incorporates motion planning principles to model rational agent behavior. Methods such as goal-anchor-based prediction play a pivotal role by estimating potential destinations and generating trajectories aligned with those goals, ensuring more interpretable and goal-driven outputs. Additionally, the integration of VLMs into prediction pipelines offers an exciting frontier, providing common-sense knowledge and contextual understanding to enhance decision-making in uncertain conditions. However, challenges like latency and scalability of these models require further investigation. Furthermore, the study underscores the importance of designing systems that generalize across diverse datasets and scenarios, addressing the existing gap in cross-dataset evaluations and qualitative assessments.
We have addressed the posed research questions, providing a comprehensive summary of the state of the art while identifying potential future research directions. Despite significant advancements over the past two decades, trajectory prediction remains an open and challenging area of research. The field has evolved from early physics-based models to Graph Transformers. The shift toward combining domain knowledge with data-driven methods, such as physics-based hybrid models, represents a crucial step forward in balancing interpretability with predictive capacity. This review not only captures the evolution of trajectory prediction methodologies but also sheds light on key research gaps, such as robust evaluation under real-world conditions and generalization to out-of-distribution scenarios. We believe this work serves as a valuable resource for researchers and practitioners, fostering innovation and inspiring further exploration in trajectory prediction domain for autonomous driving.

### v3

@@ 9.3 On Research Question 3
Recent research trends in the prediction domain focus on enhancing the robustness and generalization of prediction models. As discussed in Section 5, quantifying and mitigating uncertainty at every stage of autonomous stack is an active research area. One approach to mitigate aleatoric uncertainty is through collaborative perception, presented in Section 8, which compensates for individual perception system limitations through constructing collaborative holistic view of the environment. Planning-conditioned trajectory prediction methods, presented in Section 6, represent a promising direction by incorporating motion planning principles into prediction frameworks. By modeling planned trajectories and potential destinations, planning-conditioned approaches provide interpretable and constraint-compliant predictions, particularly in interactive and complex scenarios.
@@ 9.4 Future Directions
    Systematic Evaluation and Benchmarking.
    A major gap in motion prediction research is the lack of standardized, cross-dataset evaluation protocols. Most models are assessed using ADE and FDE on isolated datasets, each with its own format, API, and leaderboard, making fair comparison difficult. However, trajectory prediction is fundamentally about learning motion patterns, and a model trained on one dataset should ideally transfer to others under a unified evaluation framework. While some benchmarks [665, 666] support large-scale cross-dataset training, such efforts remain limited in scope. The main limitation is a unified parametrization of traffic scenarios [666] that captures interpretable factors such as road geometry, interaction density, and rare maneuvers. Without this, it is difficult to isolate failure cases, evaluate the generalization capacity of models, or identify where contextual inputs like maps or multi-agent cues are truly beneficial. A unified framework for scenario characterization, motion pattern clustering, and rare behavior sampling is needed to support robust, transferable, and reproducible model evaluation.
  4. 4.
@@
    For integration of perception and prediction, the review briefly discussed both the detect-track-predict and joint learning paradigms. The most common paradigm, detect-track-predict, propagates detection and tracking errors into the prediction stage. However, most prediction models assume clean, error-free trajectories during inference, which is far from real deployment conditions. A key research gap lies in evaluating model robustness under noisy perception and developing training strategies that allow models to generalize to imperfect inputs. Additionally, prediction-informed perception [667, 668], where future trajectories are used to guide perception, and prediction-informed tracking [669], where forecasts improve association under occlusion, also require further exploration. These approaches represent early steps toward more integrated perception–prediction pipelines but remain underexplored.
  5. 5.
@@
    Handling out-of-distribution scenarios remains an active research area. One promising direction involves a dual-process strategy: using predefined, optimized algorithms in familiar scenarios while leveraging the generalized reasoning capabilities of Vision-Language Models (VLMs) in unfamiliar or uncertain situations. Another approach focuses on simulation-to-reality (Sim2Real) adaptation. Sim2Real techniques involve generating synthetic scenes that mimic real-world conditions, enabling models to learn in controllable environments where varying levels of diversity can be introduced before deployment. Additionally, continual learning during trial deployments or pre-deployment phases can help tailor models to local conditions. A small dataset specific to the target city may be collected to fine-tune the model while preserving previously learned knowledge. A set of works in the literature explore this direction [671, 672, 149] and continues to be an active area of research. A detailed overview on this topic is provided in [152]; we refer the reader to the generalization section before pursuing further research in this direction.
  7. 7.
@@
    Autonomous vehicles must predict trajectories for diverse agent categories, including vehicles of various sizes (e.g., trucks, vans, sports cars), pedestrians, bicycles, and motorcycles. Each agent type exhibits distinct kinematic properties (except pedestrians), typical speeds, and acceleration patterns, which significantly influence their motion behavior. Vehicle trajectories are constrained by factors such as steering limits and acceleration bounds, whereas pedestrian motion is more unconstrained and inherently stochastic. A key challenge is developing a unified prediction framework capable of accommodating these heterogeneous characteristics. One approach involves using a single model trained on large, diverse datasets, with agent category labels and relevant priors provided as inputs, or following a hierarchical design where the model first classifies the agent type and then delegates prediction to category-specific modules. This setup requires carefully curated training data to ensure balanced learning across categories. Alternatively, using category-specific predictors can improve specialization but increases inference time and system complexity.
  8. 8.
@@
    Trajectory prediction models find integration in both fully autonomous vehicles and modern advanced driver-assistance systems (ADAS), contributing significantly to road safety. In full-stack autonomy, prediction modules inform planning about potential collisions. Automotive companies like Waymo and Cruise integrate multi-agent forecasting in their planning pipelines [^9] https://waymo.com/research/motionlm/. In ADAS [^10] https://www.mobileye.com/blog/ prediction models help infer lane-change intent and pedestrian movement, enabling early warnings that give human drivers more time to maneuver and avoid collisions. Another application involves integrating prediction with emergency braking systems, as seen in Mercedes-Benz vehicles, which use predictive cues to autonomously stop when drivers respond too late [^11] https://group.mercedes-benz.com/innovation/product-innovation/technology/pedestrian-emergency-braking-system.html, contributing to enhanced vehicle safety [684]. Future research should continue on integrating prediction models in traffic management systems to enable proactive safety interventions.
@@ 10 Conclusion
Beyond the taxonomy and evaluation, this study discussed the pressing challenges and evolving trends in the field. Sections on robustness and generalizability emphasize the need for uncertainty quantification and mitigation, while collaborative perception is discussed as a promising solution for addressing limitation of individual perception such as occlusions, sensor noise, and restricted sensing capabilities. The review also explored planning-conditioned trajectory prediction, which incorporates motion planning principles to model rational agent behavior. Methods such as goal-anchor-based prediction play a pivotal role by estimating potential destinations and generating trajectories aligned with those goals, ensuring more interpretable and goal-driven outputs. Additionally, the integration of VLMs into prediction pipelines offers an exciting frontier, providing common-sense knowledge and contextual understanding to enhance decision-making in uncertain conditions. However, challenges like latency and scalability of these models require further investigation. Furthermore, the study underscores the importance of designing systems that generalize across diverse datasets and scenarios, addressing the existing gap in cross-dataset evaluations and qualitative assessments.
We have addressed the posed research questions, providing a comprehensive summary of the state of the art while identifying potential future research directions. Despite significant advancements over the past two decades, trajectory prediction remains an open and challenging area of research. The field has evolved from early physics-based models to Graph Transformers. The shift toward combining domain knowledge with data-driven methods, such as physics-based hybrid models, represents a step forward in balancing interpretability with predictive capacity. This review not only presented the evolution of trajectory prediction methodologies but also shed the light on key research gaps and directions for further exploration in trajectory forecasting for autonomous driving.